OpenStudy (anonymous):

Tickets for a football game cost $8 and $12. If 480 tickets were sold for total reciepts of $4620, how many of each type of ticket were sold?

OpenStudy (anonymous):

So set up two equations: Ticket A (8 dollar) and ticket B (12 dollars) Then set up two equations like so: A+B = 480 & 8A + 12B = 4620 Now solve this as a system of equations.

OpenStudy (anonymous):

sorry im dumb how do i solve system of equations? lol

OpenStudy (anonymous):

well, set the first equation to A = 480 -B Now put this in the second equation to do: 8(480-B) + 12B = 4620 And solve for B, once you have B, put it in the first one and solve for A.

OpenStudy (anonymous):

THANKS ZELDA!

OpenStudy (anonymous):

No problem, remember to mark it best response though.